OpenStudy (anonymous):

One more! The denominator of a fraction is 1 more than the numerator. If the numerator is decreased by 1, the value of the resulting fraction is 3/4. What is the original fraction?

OpenStudy (anonymous):

Let the fraction be x/y.Now here given that y=x+1 and also (x-1)/y=3/4.So there you go two equations.Now solving

OpenStudy (diyadiya):

Pavan has given you the two equations Let me write it down again Let the fraction be \( \frac{x}{y}\) Denominator y= x+1 If the numerator is decreased by 1, the value of the resulting fraction is 3/4. \[\frac{(x-1)}{y}= \frac{3}{4}\]

OpenStudy (diyadiya):

\[y=x+1 \tag{1}\]\[\frac{x-1}{y}= \frac{3}{4}\tag{2}\] Can ypu simplify the second equation ?

OpenStudy (anonymous):

Hmmmmm maybe get rid of the fraction on the left, so it looks like this? x-1=3/4y? (not 4y but 3/4 times y?)

OpenStudy (diyadiya):

Yes, But just try to cross multiply ,it'll look simpler :)

OpenStudy (anonymous):

If we xmultiply, isn't it 4(x-1)=3y?

OpenStudy (anonymous):

Oh then plug in y=x+1 SOLVING!

OpenStudy (anonymous):

Fraction: 7/8 Diyadiya thank you so much!

OpenStudy (diyadiya):

Glad to Help :D